OpenStudy (anonymous):

3x - 9y = 3 and 6x - 3y = -24. Solve using two different methods and explain which method you found to be more efficient.

OpenStudy (anonymous):

Hopefully this webpage will help you understand this topic better: http://www.mathsteacher.com.au/year10/ch04_simultaneous/02_substitution_method/subs.htm

hero (hero):

3x - 9y = 3 6x - 3y = -24 The first thing you want to do is simply both equations to this: x - 3y = 1 2x - y = -8 It's easier to work with this way.

OpenStudy (anonymous):

How did you get 1 and -8 though?

hero (hero):

Divide both sides of the first equation by 3 Divide both sides of the second equation by 3

hero (hero):

Actually, if you only divide the first equation by three, then you'll have: x - 3y = 1 6x - 3y = -24 Then you can simply subtract the second equation from the first to eliminate the y variable.

hero (hero):

So in this case, I'd say elimination might be the easier route since it would involve less steps.

OpenStudy (anonymous):

That's true, so would x - 3y = 1 and 6x - 3y = -24 be the answer to the equations?

hero (hero):

No because you have to find the (x,y) point. So you have to find the value of x and y. Remember those equations are actually two lines that intersect at a point.